OpenStudy (sambam):

Rationalize the denominator: Please show your work so I can learn how to do it.

OpenStudy (sambam):

\[\frac{ 3\sqrt{2} }{ 2\sqrt{3}-5\sqrt{2} }\]

OpenStudy (sambam):

The bottom is in hypothesis.

OpenStudy (luigi0210):

multiply by the conjugate

OpenStudy (sambam):

how do you do that?

OpenStudy (luigi0210):

It's like multiplying by the opposite

OpenStudy (sambam):

could you explain.

OpenStudy (luigi0210):

one min

OpenStudy (luigi0210):

I've attached an answer key

OpenStudy (luigi0210):

http://www.mathsisfun.com/algebra/conjugate.html This should help with understanding conjugates